# https://www.leeds.ac.uk/

### v1

@@ Abstract
Non-prehensile manipulation in high-dimensional systems is challenging for a variety of reasons, one of the main reasons is the computationally long planning times that come with a large state space. Trajectory optimisation algorithms have proved their utility in a wide variety of tasks, but, like most methods struggle scaling to the high dimensional systems ubiquitous to non-prehensile manipulation in clutter as well as deformable object manipulation. We reason that, during manipulation, different degrees of freedom will become more or less important to the task over time as the system evolves. We leverage this idea to reduce the number of degrees of freedom considered in a trajectory optimisation problem, to reduce planning times. This idea is particularly relevant in the context of model predictive control (MPC) where the cost landscape of the optimisation problem is constantly evolving. We provide simulation results under asynchronous MPC and show our methods are capable of achieving better overall performance due to the decreased policy lag whilst still being able to optimise trajectories effectively.
@@ 1 Introduction
Figure 1: A sequence of snapshots showing an example MPC trajectory generated by our method. The task is to push the green cylinder to a goal region (the green transparent cylindrical region) through moderate clutter. The full number of DoFs in this system is 55. Our method identifies the relevant DoFs of this system at different times during execution and performs trajectory optimisation using this reduced state. Objects with stronger shades of red have more DoFs in the state vector at that point during execution: If an object is dark red, all of its six DoFs are considered; if an object is white, none of its six DoFs are considered during trajectory optimisation.
@@
Fig. 1 shows a trajectory generated by our method for an example task. In this task, the goal is to move a target object (green cylinder) to a goal region (green silhouette) whilst minimally disturbing a set of distractor objects from their initial positions. This task is motivated by the idea of a robot trying to pack an item into a tightly packed space (like a shelf or fridge) whilst minimally disturbing the environment. Traditionally, during trajectory optimisation in such a scene, all six DoFs of all movable rigid objects (which contribute twelve state vector elements, the positional and velocity element of each DoF) are considered in the system state, in addition to the robot DoFs. Instead, we propose a method that can identify the relevant DoFs at different times during the task, and perform trajectory optimisation only with the reduced system state.
@@
We make the following contributions in this paper [^1] Data, code and video can be found here:
@@ 2 Related work
Different methods for dimensionality reduction in robotics have been explored. For instance, Rapidly-exploring Random Trees (RRT) struggle to find valid motion plans as the dimensionality of the search space increases. Zheng et al. [31] consider reducing the dimensionality of the sample state space to reduce the time taken to find an optimal solution. Similarly, Vernaza et al. [27, 28] consider locomotion problems for abstract robots that have a large number of DoFs that need to navigate around a cluttered space to some desired configurations whilst avoiding collisions. A different form of dimensionality reduction is considered by Hogan and Rodriguez [7], they formulate a direct transcription trajectory optimisation problem for a 2D push-slider system and consider the exponential nature of switching contact modes over a long trajectory horizon. To limit the exponential explosion, they limit the number of times the contact mode can switch over a trajectory to make the problem computationally tractable.
@@
Ciorcarlie et al. [6] consider reducing the dimensionality of grasping problems by restricting the number of grasps to a set of key “eigengrasps”. They show that this reduced action space is still capable grasping a wide variety of complex objects. Jin et al. [9] consider learning a reduced-order hybrid model for complex three finger manipulation, enabling real-time closed loop MPC. Chen et al. [5] consider the problem of learning a reduced order model for bipedal locomotion. In their work they formulate a set of tasks for a bipedal robot to perform (walking on level ground, walking uphill etc) for which they want to compute a reduced order model offline that can be used for all their specified tasks. They manually specify a feature vector (a lower dimensional combination of features from the full-order model) and learn a set of linear weights for this feature vector offline by performing trajectory optimisation using the current reduced order model. The linear weights are adjusted via stochastic gradient descent.
These learning methods leverage prior distributions of the tasks to learn a reduced model. However, in robotic object manipulation, often the number and types of objects are unknown in advance, and each problem may require a different reduced model. We propose a method that can dynamically reduce the system state over time during MPC execution. At any given point, we use information from the trajectory optimiser to identify the most relevant DoFs for the current task configuration and optimise in this reduced state space. Our approach does not assume any prior task distribution, and we evaluate its effectiveness for both rigid and deformable object manipulation in cluttered environments.
@@ 4.1 Optimise
where $A^{\mathcal{C}}_{t}=\delta f(\textbf{x}^{\mathcal{C}}_{t},\textbf{u}_{t})/\delta\textbf{x}^{\mathcal{C}}_{t}$ and $B_{t}=\delta f(\textbf{x}^{\mathcal{C}}_{t},\textbf{u}_{t})/\delta\textbf{u}_{t}$. iLQR also requires a first and second order approximation of the cost derivatives with respect to the state and control vector $(l^{\mathcal{C}}_{\textbf{x}},l^{\mathcal{C}}_{\textbf{xx}},l_{\textbf{u}},l_{\textbf{uu}})$.
@@ 5.1 Task definition
5.1.1 Clutter
The aim of this task is to push a green cylinder to a target goal region (shown by a green silhouette) whilst minimally disturbing a set of distractor objects (example in Fig. 1). The task hyper-parameters were as follows; $T=80$, $\Delta t=0.004$, $Y=2000$, Slowdown factor = $1$.
5.1.2 Soft
The aim of this task is to push a soft body. This task aimed to push a high dimensional red soft body to a goal location (example in Fig. 2). The task hyper-parameters were as follows; $T=50$, $\Delta t=0.004$, $Y=1000$, Slowdown factor = $3$.
5.1.3 Soft rigid
This task aimed to push a green cylinder to a target location (green silhouette) but there is a high dimensional soft body in between the robot end-effector and the goal object. This means that the optimiser needs to reason about the dynamics between the soft body and the rigid body to achieve the task (example in Fig. 3). The task hyper-parameters were as follows; $T=100$, $\Delta t=0.004$, $Y=1000$, Slowdown factor = $5$.
@@ 5.2 Asynchronous MPC results
iLQR-Baseline MPC cost $1.00\pm 0.034$ $1.00\pm 0.077$ $1.00\pm 0.096$
Opt time (ms) $485.71\pm 5.16$ $620.97\pm 5.92$ $4322.39\pm 366.66$
@@
iLQR-Rand MPC cost $0.94\pm 0.033$ $1.14\pm 0.103$ $0.74\pm 0.070$
@@
iLQR-Naive MPC cost $0.89\pm 0.035$ $0.92\pm 0.066$ $0.38\pm 0.063$
@@
iLQR-SVR-SVD MPC cost $0.86\pm 0.033$ $0.90\pm 0.063$ $0.88\pm 0.070$
Opt time (ms) $335.16\pm 13.6$ $367.18\pm 49.60$ $2320.80\pm 392.85$
$\theta=$10, $\rho=$1 Num DoFs $31.85\pm 1.12$ $72.55\pm 9.37$ $85.19\pm 18.33$
iLQR-SVR-SVD MPC cost $0.91\pm 0.036$ $0.91\pm 0.066$ $0.66\pm 0.063$
@@
$\theta=$10, $\rho=$500 Num DoFs $21.06\pm 0.33$ $23.32\pm 0.77$ $24.72\pm 0.15$
iLQR-SVR-SVD MPC cost $0.91\pm 0.036$ $0.90\pm 0.066$ $0.89\pm 0.066$
Opt time (ms) $271.21\pm 11.41$ $251.90\pm 36.33$ $1296.29\pm 159.46$
$\theta=$5, $\rho=$1 Num DoFs $24.38\pm 1.00$ $49.59\pm 6.97$ $38.69\pm 7.50$
iLQR-SVR-Sum MPC cost $0.86\pm 0.032$ $0.94\pm 0.074$ $0.77\pm 0.074$
Opt time (ms) $287.05\pm 10.91$ $136.18\pm 9.76$ $1130.83\pm 48.91$
$\theta=$10, $\rho=$1 Num DoFs $29.82\pm 1.01$ $26.68\pm 1.86$ $29.10\pm 2.53$
iLQR-SVR-Sum MPC cost $0.87\pm 0.029$ $1.00\pm 0.081$ $0.70\pm 0.074$
Opt time (ms) $192.59\pm 2.08$ $101.15\pm 4.73$ $1130.83\pm 48.91$
$\theta=$10, $\rho=$500 Num DoFs $18.69\pm 0.10$ $18.20\pm 0.10$ $29.10\pm 2.53$
iLQR-SVR-Sum MPC cost $0.94\pm 0.042$ $0.97\pm 0.077$ $0.75\pm 0.066$
Opt time (ms) $218.45\pm 8.38$ $114.74\pm 12.30$ $935.45\pm 36.10$
$\theta=$5, $\rho=$1 Num DoFs $21.91\pm 0.86$ $19.94\pm 1.62$ $20.21\pm 0.92$
@@ 6 Discussion
Whilst all our methods improve the performance for all three task compared to iLQR-Baseline, our dynamic methods only have marginal gains when compared to iLQR-Naive. We believe however that the ideas presented in this work are an important step to making high-dimensional MPC possible on a wide variety of manipulation tasks.

### v2

# Online state vector reduction during model predictive control with gradient-based trajectory optimisation

@@ Abstract
Non-prehensile manipulation in high-dimensional systems is challenging for a variety of reasons. One of the main reasons is the computationally long planning times that come with a large state space. Trajectory optimisation algorithms have proved their utility in a wide variety of tasks, but, like most methods struggle scaling to the high dimensional systems ubiquitous to non-prehensile manipulation in clutter as well as deformable object manipulation. We reason that, during manipulation, different degrees of freedom will become more or less important to the task over time as the system evolves. We leverage this idea to reduce the number of degrees of freedom considered in a trajectory optimisation problem, to reduce planning times. This idea is particularly relevant in the context of model predictive control (MPC) where the cost landscape of the optimisation problem is constantly evolving. We provide simulation results under asynchronous MPC and show our methods are capable of achieving better overall performance due to the decreased policy lag whilst still being able to optimise trajectories effectively.
@@ 1 Introduction
Figure 1: A sequence of snapshots showing an example MPC trajectory generated by our method. The task is to push the green cylinder to a goal region (the green transparent cylindrical region) whilst minimally disturbing some clutter objects. The full number of DoFs in this system is 55. Our method identifies the relevant DoFs of this system at different times during execution and performs trajectory optimisation using this reduced state. Objects with stronger shades of red have more DoFs in the state vector at that point during execution: If an object is dark red, all of its six DoFs are considered; if an object is white, none of its six DoFs are considered during trajectory optimisation.
@@
Fig. 1 shows a trajectory generated by our method for a non-prehensile manipulation task in clutter. Traditionally, during trajectory optimisation in such a scene, all six DoFs of all movable rigid objects (which contribute twelve state vector elements, the positional and velocity element of each DoF) are considered in the system state, in addition to the robot DoFs. Instead, we propose a method that can identify the relevant DoFs at different times during the task, and perform trajectory optimisation only with the reduced system state.
@@
We make the following contributions in this paper111 Code and data: here:
@@ 2 Related work
Different methods for dimensionality reduction in robotics have been explored. For instance, Rapidly-exploring Random Trees (RRT) struggles to find valid motion plans as the dimensionality of the search space increases. Zheng et al. [32] consider reducing the dimensionality of the sample state space to reduce the time taken to find an optimal solution. Similarly, Vernaza et al. [28, 29] consider locomotion problems for abstract robots that have a large number of DoFs that need to navigate around a cluttered space to some desired configurations whilst avoiding collisions. A different form of dimensionality reduction is considered by Hogan and Rodriguez [7], they formulate a direct transcription trajectory optimisation problem for a 2D push-slider system and consider the exponential nature of switching contact modes over a long trajectory horizon. To limit the exponential explosion, they limit the number of times the contact mode can switch over a trajectory to make the problem computationally tractable.
@@
Notably, Sharma and Chakravorty [25] propose a reduced order iLQR implementation similar to ours, where the roll-outs are performed using the full system dynamics. They evaluate their methods for optimisation until convergence for given high-dimensional PDEs. Our work focuses on robotic manipulation problems where analytical formulations of the PDEs are not readily available. We also focus on dynamically changing the size of the state vector used during online execution.
Ciorcarlie et al. [6] consider reducing the dimensionality of grasping problems by restricting the number of grasps to a set of key “eigengrasps”. They show that this reduced action space is still capable grasping a wide variety of complex objects. Jin et al. [9] consider learning a reduced-order hybrid model for complex three finger manipulation, enabling real-time closed loop MPC. Chen et al. [5] consider the problem of learning a reduced order model for bipedal locomotion. In their work, they formulate a set of tasks for a bipedal robot to perform (walking on level ground, walking uphill etc) for which they want to compute a reduced order model offline that can be used for all their specified tasks. They manually specify a feature vector (a lower dimensional combination of features from the full-order model) and learn a set of linear weights for this feature vector offline by performing trajectory optimisation using the current reduced order model. The linear weights are adjusted via stochastic gradient descent.
These learning methods leverage prior distributions of the tasks to learn a reduced model. However, in robotic object manipulation, often the number and types of objects are unknown in advance, and each problem may require a different reduced model. We propose a method that can dynamically change the system state over time during MPC execution. At any given point, we use information from the trajectory optimiser to identify the most relevant DoFs for the current task configuration and optimise in this reduced state space. Our approach does not assume any prior task distribution, and we evaluate its effectiveness for both rigid and deformable object manipulation in cluttered environments.
@@ 4.1 Optimise
where $A^{\mathcal{C}}_{t}=\delta f(\textbf{x}^{\mathcal{C}}_{t},\textbf{u}_{t})/\delta\textbf{x}^{\mathcal{C}}_{t}$ and $B_{t}^{\mathcal{C}}=\delta f(\textbf{x}^{\mathcal{C}}_{t},\textbf{u}_{t})/\delta\textbf{u}_{t}$. iLQR also requires a first and second order approximation of the cost derivatives with respect to the state and control vector $(l^{\mathcal{C}}_{\textbf{x}},l^{\mathcal{C}}_{\textbf{xx}},l_{\textbf{u}},l_{\textbf{uu}})$.
@@ 5.1 Task definition
Clutter task: The aim of this task is to push a green cylinder to a target goal region (shown by a green silhouette) whilst minimally disturbing a set of distractor objects (example in Fig. 1). The task hyper-parameters were as follows; $T=80$, $\Delta t=0.004$, $Y=2000$, Slowdown factor = $1$.
Soft task: The aim of this task is to push a soft body. This task aimed to push a high dimensional red soft body to a goal location (example in Fig. 2). The task hyper-parameters were as follows; $T=50$, $\Delta t=0.004$, $Y=1000$, Slowdown factor = $3$.
Soft rigid task: This task aimed to push a green cylinder to a target location (green silhouette) but there is a high dimensional soft body in between the robot end-effector and the goal object. This means that the optimiser needs to reason about the dynamics between the soft body and the rigid body to achieve the task (example in Fig. 3). The task hyper-parameters were as follows; $T=100$, $\Delta t=0.004$, $Y=1000$, Slowdown factor = $5$.
@@ 5.2 Asynchronous MPC results
iLQR-Baseline MPC cost $1.00\pm 0.034$ $1.00\pm 0.080$ $1.00\pm 0.096$
Opt time (ms) $485.71\pm 5.16$ $620.97\pm 7.35$ $4322.39\pm 366.65$
@@
iLQR-Rand MPC cost $0.94\pm 0.033$ $1.14\pm 0.104$ $0.74\pm 0.071$
@@
iLQR-Naive MPC cost $0.89\pm 0.035$ $0.92\pm 0.067$ $0.38\pm 0.066$
@@
iLQR-SVR-SVD MPC cost $0.86\pm 0.033$ $0.90\pm 0.065$ $0.88\pm 0.071$
Opt time (ms) $335.16\pm 13.61$ $367.18\pm 49.60$ $2320.80\pm 392.91$
$\theta=$10, $\rho=$1 Num DoFs $31.85\pm 1.12$ $72.55\pm 9.37$ $85.19\pm 18.32$
iLQR-SVR-SVD MPC cost $0.91\pm 0.036$ $0.91\pm 0.069$ $0.66\pm 0.063$
@@
$\theta=$10, $\rho=$500 Num DoFs $21.06\pm 0.33$ $23.32\pm 0.77$ $24.72\pm 0.17$
iLQR-SVR-SVD MPC cost $0.91\pm 0.036$ $0.90\pm 0.061$ $0.89\pm 0.068$
Opt time (ms) $271.21\pm 11.41$ $251.90\pm 38.30$ $1296.29\pm 184.13$
$\theta=$5, $\rho=$1 Num DoFs $24.38\pm 1.00$ $49.59\pm 7.34$ $38.69\pm 8.66$
iLQR-SVR-Sum MPC cost $0.86\pm 0.032$ $0.94\pm 0.073$ $0.77\pm 0.074$
Opt time (ms) $287.05\pm 10.91$ $136.18\pm 10.92$ $1130.83\pm 54.68$
$\theta=$10, $\rho=$1 Num DoFs $29.82\pm 1.01$ $26.68\pm 2.08$ $29.10\pm 2.92$
iLQR-SVR-Sum MPC cost $0.87\pm 0.029$ $1.00\pm 0.081$ $0.70\pm 0.071$
Opt time (ms) $192.59\pm 2.08$ $101.15\pm 4.73$ $997.71\pm 32.55$
$\theta=$10, $\rho=$500 Num DoFs $18.69\pm 0.10$ $18.20\pm 0.11$ $24.09\pm 0.46$
iLQR-SVR-Sum MPC cost $0.94\pm 0.042$ $0.97\pm 0.079$ $0.75\pm 0.066$
Opt time (ms) $218.45\pm 8.38$ $114.74\pm 13.76$ $935.45\pm 37.03$
$\theta=$5, $\rho=$1 Num DoFs $21.91\pm 0.86$ $19.94\pm 1.87$ $20.21\pm 0.91$
@@ 6 Discussion
Acknowledgments
This research has received funding from the UK Engineering and Physical Sciences Research Council under the grants EP/V052659/1 and 2596473. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising.